FEL Upgrade Project Weekly Brief
February 26-March 2, 2001



Users of the FEL had a successful week at materials processing including studies of laser nitriding and laser hardening.
Peter Schaaf from the University of Goettingen was given most of the week to continue his investigations of laser nitriding.
He appears to have successfully nitrided Fe, SS, Al, Ti and Si samples  using a simple x-y scanner with a flowing nitrogen
purge. We also delivered light for amorphous metal layer production on iron samples  for Greg Kessel from Virginia


On Thursday March 1, John Eric from the Air Force Research Lab visited Jefferson Lab for a briefing and a working
meeting on the contractual arrangements for the kW UV project.  A statement of work and management plan has been
submitted that is acceptable to AFRL and all parties are working toward an April start of the project.

On Friday March 2 we were visited by 8 members of the DoD  Joint Technology Office and 4 visitors from Lincoln Lab
and NRL.  The JTO has recently funded a study using the FEL for short pulse materials damage effects.  After general
discussions it was agreed that NRL would discuss
potential collaboration with JLab on providing optical and plasma diagnostics for some of the planned studies

The project monthly report for January was released this week to the DOE and Navy contract monitors.

WBS 3 (Beam Physics):

Staff participated in VBA/Excel training in support of machine modeling work.

WBS 4 (Injector):

Refinement of the internal cesiation scheme was started.  The first cut of the gun chamber drawings will be done early next

Gun HVPS - A quote of $1500 was received for reusable shipping containers for the Gun HVPS.  Glassman needs our
spare HVPS make the upgrade to it.

WBS 6 (RF):

Zone 3 - Contract with Hippotronics for the Cathode Power Supply is ready for signatures.  Many parts are still missing,
but are on order.

Zone 4 - The software to test the zone is to be completed 3/09/01.  A hardware punch list is being worked and is
expected to be completed by 3/16/01 with the exception of a missing circulator.  The circulator failed due to low coolant
flow and we have no spares on site.

Injector HVPS - PSC is reviewing their offer.  A decision and price is due next Friday.  They are reluctant to accept this

Injector 100 kW Klystron - CPI is having late delivery problems with their vendors.  They will be late in delivering the
100 kW klystrons, but this will have no impact on the project.

WBS 8 (I&C):

The new printed circuit board for OBPMs were received and tested.  Still some more work to be done on this.  The
original detectors are 4 elements 1mm x 1mm separated by a 75 micron gap, lithium tantalate from Eltec. These had gold
bonded wires with anchor pads < 100 microns in diameter, located near the outside edge of the detectors, the next batch
that we purchased had pads > 200 microns in diameter that were located near the center of the 1mm x 1mm element.
When the FEL beam would fall upon the pad the energy would be reflected, indicating an error in position. Two
Molectron Q5's were purchased by Michelle, these are in and being evaluated. These are also lithium tantalate but are
5mm in diameter and split into four segments. The new picomotor controller chassis, which will control the transport
mirrors, has been tested on the bench and is ready for installation.<?xml:namespace prefix = o ns
="urn:schemas-microsoft-com:office:office" />

S-Video has been added to the television in the conference area so it can be used for computer presentations, etc.
Upgrades continue in the user labs for motor control and video. The first of eight user power supplies is nearly complete,
the balance will be fabricated commercially.  These will supply power to cameras and optical instrumentation.  Cameras
were installed for the user in Lab 3 hutch.

Discussions continue for a commercial trim power supply. Efforts are centered on the control philosophy and architecture.

The parts continue to arrive for the vacuum control system and power supplies. Fabrication shops are being queried for
the ion pump power supplies assembly.

Two outstanding drawings were signed off and submitted to Document Control.  Two others were submitted to EECAD
on 3/01 for the final up-dates and should be signed off soon.  A set of drawings has been made for the wiring of the
OBPM pico motor control chassis. It will be handed over to EECAD at the next available opportunity. Many others are
still being up-dated.

WBS 9 (Transport):

Injector Dipoles (DU/DV)
o  DULY Research obtained a field that was within the flatness specification using their 3D model
     (in RADIA) of a parallel faced dipole and mu metal.  Their results with the actual wedge model
    still have a gradient that is an order of magnitude out of specification, and (counter intuitively)
    still high on the narrow side.  They now have enough confidence in their model that they will
    make changes to investigate the source of the gradient.  Their drawing set and the drafting
    model of the GV (Small injector dipole) include poles, field clamps and coils. Work is ongoing on
    the leads and water connections.
 Optical Chicane Dipole (DW)
o  We generated the coil drawings for the DW drawing package and continue back check of the
    remaining parts.
Arc Dipoles (GY, GX, GQ)
o  AES made substantial progress in designing GY (180 degree dipole) coil's leads, water
    connections, bus arrangement, electrical cover and external leads into a single system.
o  AES worked at debugging their 3D magnetic model of the GY
QX (3.125" Quad)
o Measurement Probe: Measurements are being made on wire thickness variations to optimize
o Design Package: The QX drawing package remains in checking.
Sextupole with H/V Correctors
o  Robin Wines continued work on the magnetic model of the Sextupole. A reasonable sextupole
    was obtained and the proposed octupole generation coils were inserted on the four poles at 30
    degrees to the horizontal.  The octupole component is indeed generated, but quadrupole field is
    also generated that is an order of magnitude higher than the octupole.  We believe this may not
    be what Dave Douglas had in mind.
o The layout of the IR lattice was extended into the injection region with study of the use of the
    existing dipole girders.

WBS 10 (Wiggler):

Awaiting dispersion section.  Measurement HALL probe has been ordered.

WBS 11 (Optics):

We received a budgetary estimate on the drawing package for the deformable mirror assembly that is in line with previous
estimates.  We will now release a requisition and finish getting quotations.  More parts for the mirror test stand arrived,
namely the new gimbal mount from Newport.  This will be used in the mirror test stand, and will be seriously evaluated for
its suitability in the upgrade optical cavity.  We anticipate the vacuum vessel for the mirror test stand to be received
next week.

We received the new quadrant pyroelectric detectors from Molectron, as well as the rebuilt high power laser beam dump.

In the course of supporting operations, we supplied parts and technical assistance to our primary users, Peter Schaff and
Greg Kessel.  We assisted in reconfiguring the hutch in User Lab 3 for metal-treatment research.  We had to deal with the
aftermath of the water seal failure last week; water entrained in the table is leaking down onto the Brewster window in the
optics control room (OCR).  We erected a "chimney" around the window to divert water from the window and are
monitoring the situation.  It did not impede operations for most of the week.


  See "Highlights" above.

Privacy and Security Notice